# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v16

@@ 2.3 Statistical Taylor Expansion
Formula (2.5) provides the uncertainty distribution of an analytic function. However, in most scientific and engineering calculations, the primary interest lies not in the full result distribution but in few summary statistics of the result, such as the mean and deviation [1][2]. These simplified statistics can be obtained through a statistical Taylor expansion.
@@
$\zeta(n)\equiv$ $\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
@@
$\overline{f(x)}=$ $\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\sum_{n=0}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.9)
@@
An confidence interval [4] $\tilde{x}\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ can describe a sampling from an underlying distribution, where $0<\varrho,\kappa$ specify the bounding ranges. Formula (2.7) defines the bound moment $\zeta(n)$. As discussed later, $\kappa$ determines $\varrho$, while $\kappa$ itself is determined by both the sample size $N$ and the underlying distribution of the input. When $N\rightarrow\infty$, $\varrho,\kappa\rightarrow\infty$, so that by definition $\zeta(0)=1$, $\zeta(1)=0$, and $\zeta(2)=1$. When $N$ is bounded, $\varrho$ and $\kappa$ are also bounded, so that $\zeta(0)<1$, and $\zeta(2)<1$. The probability of $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.8). Using Formula (2.7), Formula (2.9) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
$f(x+\tilde{x},y+\tilde{y})=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\frac{f^{(m,n)}_{(x,y)}}{m!n!}\tilde{x}^{m}\tilde{y}^{n};$ (2.11)
$\overline{f(x,y)}=$ $\int\int f(x+\tilde{x},y+\tilde{y})\rho(\tilde{x},\mu_{x},\sigma_{x})\rho(\tilde{y},\mu_{y},\sigma_{y})\;d\tilde{x}d\tilde{y}$ (2.12)
$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$
@@
Under the uncorrelated uncertainty assumption, Formula (2.12) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.11), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively. Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
2.4 Bound Moment Symmetry
A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbounded distribution. To achieve mean preserving, $\kappa$ determines $\varrho$, so that $\zeta(n)$ becomes $\zeta(n,\kappa)$. Under a mean preserving bounding, Formula (2.14) and (2.15) provide the result for $x\pm y$, while Formula (2.16) and (2.17) give the result for $xy$:
@@ 2.4 Bound Moment Symmetry
When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.14) and (2.15) the convolution results for $x\pm y$ [4], and Formula (2.16) and (2.17) the corresponding results of the product distribution for $xy$ [4].
For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
@@ 2.5 One-Dimensional Examples
$e^{x+\tilde{x}}=e^{x}\sum_{n=0}^{\infty}\frac{\tilde{x}^{n}}{n!};$ (2.18)
$\frac{\overline{e^{x}}}{e^{x}}$ $=\sum_{n=0}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.19)
@@
$\log(x+\tilde{x})-\log(x)=\log(1+\frac{\tilde{x}}{x})=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j}\frac{\tilde{x}^{j}}{x^{j}};$ (2.21)
@@
$(x+\tilde{x})^{c}=x^{c}(1+\frac{\tilde{x}}{x})^{c}=x^{c}+x^{c}\sum_{n=1}^{\infty}\frac{\tilde{x}^{n}}{x^{n}}\begin{pmatrix}c\\ n\end{pmatrix};\;\;\;\begin{pmatrix}c\\ n\end{pmatrix}\equiv\frac{\prod_{j=0}^{n-1}(c-j)}{n!};$ (2.27)
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+1\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.28)
@@
The result variance in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
2.6 Low-Order Approximation
When $n<5\leq\kappa$, $\eta(n)\simeq n!!$. Under these conditions, Formula (2.20), (2.23), (2.26), and (2.29) can be simplified as Formula (2.30), (2.31), (2.6), and (2.6), respectively.
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}\simeq$ $\;\sigma^{2}+\frac{3}{2}\sigma^{4}+\frac{7}{6}\sigma^{6}+\frac{5}{8}\sigma^{8}+o((\delta x)^{10});$ (2.30)
$\delta^{2}\log(x)\simeq$ $\;P(x)^{2}+P(x)^{4}\frac{9}{8}+P(x)^{6}\frac{119}{24}+P(x)^{8}\frac{991}{32}+o(P(x)^{10});$ (2.31)
$\delta^{2}\sin(x)\simeq$ $\;\sigma^{2}\cos(x)^{2}-(\delta x)^{4}(\cos(x)^{2}\frac{3}{2}-\frac{1}{2})$
$+\sigma^{6}(\cos(x)^{2}\frac{7}{6}-\frac{1}{2})-\sigma^{8}(\cos(x)^{2}\frac{5}{8}-\frac{7}{24})+o((\delta x)^{10});$ (2.32)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}\simeq$ $\;c^{2}P(x)^{2}+\frac{3}{2}c^{2}(c-1)(c-\frac{5}{3})P(x)^{4}$
$+\frac{7}{6}c^{2}(c-1)(c-2)^{2}(c-\frac{16}{7})P(x)^{6}+o(P(x)^{8});$ (2.33)
Formula (2.34), (2.35), and (2.36) are special cases of Formula (2.6).
$\delta^{2}x^{2}\simeq$  $\;4x^{2}(\delta x)^{2}+2(\delta x)^{4};$ (2.34)
$\frac{\delta^{2}\sqrt{x}}{(\sqrt{x})^{2}}\simeq$ $\;\frac{1}{4}P(x)^{2}+\frac{7}{32}P(x)^{4}+\frac{75}{128}P(x)^{6}+o(P(x)^{8});$ (2.35)
$\frac{\delta^{2}1/x}{(1/x)^{2}}\simeq$ $\;P(x)^{2}+8P(x)^{4}+69P(x)^{6}+o(P(x)^{8});$ (2.36)
2.7 Asymptotic Behavior for Bound Moment
When $\rho(\tilde{z})>0$ and $\frac{d\rho(\tilde{z})}{d\tilde{z}}$ is bounded in $\tilde{z}\in[0,\kappa]$, Formula (2.38) and (2.39) leads to Formula (2.40). Formula (2.37) and (2.40) conclude Formula (2.41) and (2.42).
$\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$ $=\rho(\kappa)\frac{\kappa^{n+1}}{n+1}-\frac{1}{n+1}\int_{0}^{\kappa}\tilde{z}^{n+1}\frac{d\rho(\tilde{z})}{d\tilde{z}}d\tilde{z}$; (2.37)
@@ 2.7 Asymptotic Behavior for Bound Moment
$\lim_{n\rightarrow+\infty}$ $\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}=\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.40)
$|\varrho|=\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
$|\varrho|<\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(n,\kappa)=\rho(\varrho)\frac{\varrho^{n+1}}{n+1}+\rho(\kappa)\frac{\kappa^{n+1}}{n+1}\simeq\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.42)
2.7.1 Gaussian Input Uncertainty
Figure 5: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus Formula (2.45) (left y-axis), and their ratio (right y-axis).
The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [4]. This convergence occurs rapidly [29]. In digital computation, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sums of expansion terms [7][8]. Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [1][2][4].
With Formula (2.43) [Double_Factorial], for Gaussian distribution, Formula (2.44) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, while Formula (2.45) demonstrates its asymptotic behavior. Figure 5 shows the transition to Formula (2.45) with increasing order $2n$, in which the ratio is the ratio of Formula (2.45) when $\kappa=5$ to $\zeta(2n,5)$.
$\sum_{j=0}^{+\infty}\frac{\kappa^{2j+1}}{(2j+1)!!}=\sqrt{\frac{\pi}{2}}\xi(\frac{\kappa}{\sqrt{2}})e^{\frac{1}{2}\kappa^{2}};$ (2.43)
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}N(\tilde{z})d\tilde{z}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.44)
$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.45)
2.7.2 Uniform Input Uncertainty
As shown in Figure 38, rounding errors can be uniformly distributed [29].
For uniform distribution, the bounding range $\kappa$ is less than the maximal range $\sqrt{3}$. Formula (2.46) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.46)
2.8 Convergence of Variance
$\lim_{n\rightarrow+\infty}$ $\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.47)
$\delta^{2}\log(x\pm\delta x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}$
$\simeq$  $\;2\nu(\kappa)\log(2)\sum_{n=1}^{+\infty}\frac{(P(x)\kappa)^{2n}}{(2n)^{2}},\begin{cases}\text{Gaussian}:\nu(\kappa)=N(\kappa)\kappa\\ \text{Uniform}:\nu(\kappa)=1,\;\;\;\kappa=\sqrt{3}\end{cases};$ (2.48)
$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.49)
For input distribution without bounding, Formula (2.3) may not converge, e.g., when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$. Thus, it is necessary that the input uncertainty is bounded. Formula (2.47) shows that the asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
Formula (2.20) for $e^{x\pm\delta x}$ and Formula (2.26) for $\sin(x\pm\delta x)$ both converge unconditionally. However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can become negative for large $\delta x$, which imposes an upper-bound constraint on the input $\delta x$.
Formula (2.23) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) as $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$.
Formula (2.29) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.49) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ . This expression converges when $P(x)\lesssim 1/\kappa$ although the precis upper bound for $P(x)$ varies with $c$.
@@ 2.9 Statistical Bounding
When the underlying distribution is Normal, Formula (2.50) and (2.51) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 6 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 7 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$). Figure 7 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$. Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 6 should be used as $\kappa$ in Formula (2.7).
When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to bounding leakage $\epsilon(N)$. Figure 7 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically. The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.7).
[…]
@@ 2.10 Ideal Statistics
When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}\rightarrow\sigma$ in Formula (2.52). Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
  • When the underlying uncertainty distribution is Gaussian, Figure 8 shows that the resulting variance $\delta^{2}f$ for a selected functions rise with NN until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 50$. When $N=20$ and $\kappa_{s}=5$, $\kappa=3.8$ according to Figure 6. Therefore it is sufficient to compute the stable variances $\widehat{\delta^{2}}f$ with $\zeta(n,5)$.
  • When the underlying uncertainty distribution is Uniform, Figure 9 shows the same trend, however the stable variances are reached only when $N>10^{4}$. The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.46).
Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input.
In both Figure 8 and 9, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding ideal variances with exactly the same trend. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$. Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution. Define the ideal ratio to be zero when the sample count $N$ is $1$. When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.53) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
An ideal ratio also applies to the corresponding uncertainty bias. The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 8 and 9, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
In statistical Taylor expansion, every calculation should be carried out twice:
  1. 1.
[…]
  2. 2.
    With the real $\zeta(n,\kappa)$ of each input given the sample count and input uncertainty distribution, to obtain the ideal ratio to quantify the reliability of the ideal variance and ideal uncertainty bias as a value in the range of $[0,1)$.

### v17

@@ 2.3 Statistical Taylor Expansion
$\zeta(n)$ $\equiv\int_{\varrho}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.2)
$\zeta(1)$ $=0;$ (2.3)
Let $\tilde{z}\in[\varrho,\kappa]$ where $\varrho,\kappa$ specify the bounding ranges. Formula (2.2) defines the corresponding bound moment $\zeta(n)$, which further satisfies the mean-reverting condition of Formula (2.3) so that $\kappa$ determines $\varrho$. For any symmetric $\rho(-\tilde{z})=\rho(\tilde{z}):\varrho=-\kappa,\zeta(2n+1)=0$. The probability of $\tilde{z}\not\in[\varrho,\kappa]$ is defined as the bounding leakage $\epsilon(\kappa)=1-\zeta(0,\kappa)$.
@@
$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=f(x)+\sum_{n=1}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.5)
@@
An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.4). Formula (2.5) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
[…]
$\overline{f(x,y)}$ $=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$ (2.8)
@@
Under the uncorrelated uncertainty condition, Formula (2.8) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.7), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively. Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
With the mean reverting condition of Formula (2.3):
@@
When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formula (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
@@ 2.4 One-Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.14)
@@
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.21)
@@
The input and output in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
2.5 Convergence of Variance
Figure 3: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus $2N(\kappa)\frac{\kappa^{2n+1}}{2n+1}$ (left y-axis), and their ratio (right y-axis).
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.23)
$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.24)
$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.25)
@@ 2.5 Convergence of Variance
[…]
  • Formula (2.15) for $e^{x\pm\delta x}$ and Formula (2.20) for $\sin(x\pm\delta x)$ both converge unconditionally.
[…]
  • Formula (2.22) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.27). which converges when $P(x)\lesssim 1/\kappa$ although the precise upper bound for $P(x)$ varies with $c$.
@@ 2.6 Statistical Bounding
When the underlying distribution is Normal, Formula (2.28) and (2.29) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases, but converges to a larger stable bounding leakage $\epsilon(\kappa_{s})$.
When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to the bounding leakage $\epsilon(N)$. Figure 5 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically. The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$.
@@ 2.7 Ideal Statistics
When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}$ in Formula (2.30). Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
  • When the underlying uncertainty distribution is Gaussian and $\kappa_{s}=5$, Figure 6 shows that the resulting variances $\delta^{2}f$ for a selected functions rise with $N$ until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 50$.
  • When the underlying uncertainty distribution is Uniform, $\kappa_{s}=\sqrt{3}$, Figure 7 shows the same trend, however the stable variances $\widehat{\delta^{2}}f$ are reached only when $N\geq 10^{4}$.
The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 6 and 7 respectively, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.5) and Formula (2.3) contains $\sigma^{2}\zeta(2,\kappa)$.
In both Figure 6 and 7, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\widehat{\delta^{2}}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding stable variances with exactly the same trend in each figure. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer the confidential interval range $\kappa$ from bounding leakage $\epsilon(\kappa_{s})=1-\zeta(0,\kappa_{s})$ in Figure 4, or more precisely $1-\zeta(2,\kappa_{s})$ for $\delta^{2}f$.
Define the stable variance when the sample count $N$ is sufficiently large as the ideal variance $\widehat{\delta^{2}}f$. The choice of $\kappa_{s}$ to calculate $\widehat{\delta^{2}}f$ is defined as the ideal bounding range $\hat{\kappa}$. For Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$. When the input distribution is unbounded, the choice of $\hat{\kappa}$ is a trade-off: a larger $\hat{\kappa}$ means a smaller bounding leakage for the result but a narrower convergence range for the input. For Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$.
When the sample count $N$ is insufficient, $\widehat{\delta^{2}}f$ is calculated using $\hat{\kappa}$, to assume that $N$ is sufficient. $\delta^{2}f$ is calculated with the actual $N$. Define the ideal ratio as $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f$, which is 0 when $N=1$ and $1$ when $N$ is sufficiently large. The ideal ratio $\alpha$ quantifies the reliability of $\widehat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.31) as an example for $\alpha(x\pm y)$ .